the sum of the digit is 12. the number is 1 more than 8 times its unit digit. find the number?
Answer by dfrazzetto(277)

for a number, xy
x + y = 12
"xy" = 1 + 8y
try different combinations
93, no
84, no
75, no
66, no
57, yes because 8(7) = 56 + 1 = 57